a rectangle is 8 feet long and 6 feet wide. if each dimension is increased by the same number of feet, a new rectangle is formed whose area exceeds the area if the original rectangle by 72 square feet. find the dimensions of the new rectangle.
Answer by richwmiller(17219)

let a=area of rectangle with sides 8 ft and 6 ft
add x to each dimension
what is the formula for area of a rectangle?
formula=a
formula2=a+72=(8+x)(6+x)
what is a?
a=Length * width
substitute a below
a+72=(8+x)(6+x)
multiply out and find x